OpenStudy (anonymous):

Every day, Marco runs for 45 minutes and covers 5 miles. How long would it take him to run 8 miles? a) 9 min b) 18 min c) 48 min d) 72 min

OpenStudy (anonymous):

8/5 * 45=72

OpenStudy (anonymous):

He takes 45 mins to cover 5 miles, so if we divide both by 5 we can find how long it takes to cover 1 mile. 45/5 = 9 mins. Now we just times this by 8 to find how long it takes him to run 8 miles. 8x9=72 mins.

OpenStudy (anonymous):

Ok if he runs 5 miles in 45 minutes, then he runs a mile every 9 minutes. So 8 times 9 is 72

OpenStudy (anonymous):

45:5 = x:8 5x = 360 x = 72 minutes using ratio and proportion

OpenStudy (anonymous):

Thanks ALLL of you, you ALL deserve medals :D

OpenStudy (anonymous):

No problem, Have a good one!

OpenStudy (shane_b):

\[Speed=\frac{distance}{time}\]Just rearrange it as necessary and you can answer any of these times of questions.

OpenStudy (anonymous):

18 min, sngles of triangles sum up to 180.

OpenStudy (shane_b):

*TYPEs of questions :)

hero (hero):

Alternatively, set up the following proportion \(\huge\frac{45 \space \text{min}}{x \space \text{min}} = \frac{5 \space \text{mi}}{8 \space \text{mi}}\) Then cancel min and mi to get \(\huge\frac{45}{x} = \frac{5}{8}\) Then cross multiply to get \(\huge 5x = 45(8)\) Then continue solving for x